Solve:


x²y²(2ydx + xdy) - (5ydx + 7xdy) =0

1
Expert's answer
2022-01-25T17:02:42-0500

We write the equation as



"2x^2y^3dx+x^3y^2dy-5ydx-7xdy=0"

Let's select the total differential x^3y^3



"\frac{1}{3} d(x^3y^3)+x^2y^3dx-5ydx-7xdy=0."

Let's select the total differential -5xy



"\frac{1}{3} d(x^3y^3)-5d(xy)+x^2y^3dx-2xdy=0."

Consider the equation



"x^2y^3dx-2xdy=0"

get



"xdx=\frac{2dy}{y^3}."

As result



"\frac{x^2}{2}+\frac{1}{y^2}=C"

where C is constant.

Hence. the general solution of the equation is



"\frac{x^3y^3}{3}-5xy+\frac{x^2}{2}+\frac{1}{y^2}=c"